### v1

# PERC: Plan-As-Query Example Retrieval for Underrepresented Code Generation

Abstract
Code generation with large language models has shown significant promise, especially when employing retrieval-augmented generation (RAG) with few-shot examples. However, selecting effective examples that enhance generation quality remains a challenging task, particularly when the target programming language (PL) is underrepresented. In this study, we present two key findings: (1) retrieving examples whose presented algorithmic plans can be referenced for generating the desired behavior significantly improves generation accuracy, and (2) converting code into pseudocode effectively captures such algorithmic plans, enhancing retrieval quality even when the source and the target PLs are different. Based on these findings, we propose Plan-as-query Example Retrieval for few-shot prompting in Code generation (PERC), a novel framework that utilizes algorithmic plans to identify and retrieve effective examples. We validate the effectiveness of PERC through extensive experiments on the CodeContests, HumanEval and MultiPL-E benchmarks: PERC consistently outperforms the state-of-the-art RAG methods in code generation, both when the source and target programming languages match or differ, highlighting its adaptability and robustness in diverse coding environments.
PERC: Plan-As-Query Example Retrieval for
Underrepresented Code Generation
Jaeseok Yoo1 Hojae Han1 Youngwon Lee1 Jaejin Kim12 Seung-won Hwang12thanks: Corresponding author. 1Seoul National University 2Interdisciplinary Program in Artificial Intelligence, Seoul National University {jaeseok2.yoo,stovecat,ludaya,jaejin.kim,seungwonh}@snu.ac.kr
Figure 1: Two Python and Lua code snippets have different modalities but implement the same algorithmic plans. PERC uses pseudocode describing the algorithmic plans to minimize noise from modality differences. Colors in code represent equivalent steps.
1 Introduction
Code generation using large language models (LLMs) has shown significant potential, particularly when retrieval-augmented generation (RAG) with few-shot examples is employed Parvez et al. (2021); Nashid et al. (2023); Zhang et al. (2023). However, selecting effective examples to improve code generation quality remains a challenging task. This is even more difficult when the target programming language (PL) is underrepresented, as the construction of the few-shot example pool for retrieval is non-trivial.
To construct the retrieval pool for an underrepresented PL, we can transfer the retrieval pool from a high-resource PL. However, this in turn interferes with the state-of-the-art few-shot prompting approaches in code generation Nashid et al. (2023); Zhang et al. (2023), which employ code to retrieve examples. Figure 1 illustrates such an example; although the Python code on the lower left and the Lua code on the lower right follow the same algorithmic steps, both lexical-based Robertson and Zaragoza (2009) and embedding-based (Song et al., 2020) retrieval fall short in capturing their algorithmic similarity due to syntactic and structural difference (An et al., 2023).
To overcome this, we propose Plan-as-query Example Retrieval for few-shot prompting in Code generation (PERC). PERC leverages algorithmic plans such as pseudocode to retrieve examples. Converting code to pseudocode reduces syntactic noise and captures algorithmic similarity, enhancing retrieval quality across programming languages. Also, such plans can aid generation by participating in reasoning chains, further improving the generation accuracy.
We demonstrate PERC’s effectiveness in two key scenarios: First, plan-based example selection improves code generation accuracy for same-language tasks on CodeContests Li et al. (2022) and HumanEval Chen et al. (2021). Second, on MultiPL-E Cassano et al. (2023), PERC enhances code generation for underrepresented languages by leveraging data from high-resource languages.
Our contribution is three-fold:
  • We propose PERC, a novel framework of leveraging algorithmic plans for few-shot example retrieval in code generation.
  • We demonstrate that plan-based retrieval improves same-language code generation on competitive programming and general-purpose coding tasks.
  • We confirm that PERC can leverage high-resource PLs to improve code generation accuracy in underrepresented PLs.
2 Related Work
Retrieval-Augmented Code Generation
Previous works adopting RAG in code generation tasks have primarily focused on enhancing the accuracy of generated code by retrieving from the target PL pool (Parvez et al., 2021; Lu et al., 2022). More recently, CEDAR Nashid et al. (2023) retrieved few-shot examples based on code-code similarity, and RepoCoder Zhang et al. (2023) leveraged LLM-generated code snippets in target PL to expand queries, allowing for improved retrieval.
Utilizing Algorithmic Plans in Code Retrieval and Generation
Han et al. (2021) viewed pseudocode as algorithmic plan of code, to reduce the modality gap between text and code in code search task. Jiang et al. (2024) used pseudocode-based algorithmic plans for code generation through few-shot prompting, and Sun et al. (2024) used pseudocode to bridge different programming languages.
Our distinction.
We are the first to leverage algorithmic plans in retrieval-augmented code generation, which allows to retrieve effective few-shot examples by reducing lexical bias. As a result, our proposed PERC naturally adapts to source-target PLs mismatches.
Figure 2: Overview of PERC: (1) plan-based retrieval and (2) code generation with few-shot prompting. PERC retrieves examples with the most similar algorithmic plans. Yellow and blue respectively signifies the triplet of the selected few-shot example and the target problem.
3 Preliminaries
Given a natural language query $t_{q}$ describing a desired program, code generation aims to return the corresponding implementation. In few-shot example retrieval, we draw relevant text-code pairs $(t,c)$ from an example pool $\mathcal{P}$ to supplement LLM’s knowledge and guide generation.
Problem-As-Query
A baseline approach maps query $t_{q}$ and text descriptions $t$ into a shared latent space using encoder $\psi$. The top-$k$ examples are selected based on similarity $\mathrm{sim}(\cdot)$:
$E=\mathrm{topk}_{{(t,c)}\in\mathcal{P}}\ \mathrm{sim}(\psi(t_{q}),\psi(t)),$ (1)
where $E$ is the set of indices of chosen examples.
CEDAR
Nashid et al. (2023) selects examples based on code-code similarity. As the prompt lacks code for querying, we use the LLM’s predicted code $\hat{c}_{q}$ from $t_{q}$ to retrieve examples:
$E_{\textrm{CD}}=\mathrm{topk}_{(t,c)\in\mathcal{P}}\ \mathrm{sim}(\psi(\hat{c}_{q}),\psi(c)).$ (2)
RepoCoder
Zhang et al. (2023) generates code $\hat{c}_{q}$ from $t_{q}$ and expands the query, following LLM-based query expansion Wang et al. (2023). The retrieval combines both problem description and code:
$E_{\textrm{RC}}=\mathrm{topk}_{(t,c)\in\mathcal{P}}\ \mathrm{sim}(\psi(t_{q};\hat{c}_{q}),\psi(t;c)),$ (3)
where semicolon denotes concatenation.
 Benchmark CodeContests HumanEval MultiPL-E
Target PL Python Python Rust Julia Lua Ruby Go R
Data Availability High-resource High-resource High-resource $\xleftrightarrow{\hskip 28.45274pt\text{}\hskip 28.45274pt}$ Underrepresented
Few-shot Prompting Method Pass@1 (%)
w/o Examples 2.72 73.17 62.82 51.45 49.94 47.76 30.91 14.78
Random Selection 5.82 72.56 61.86 52.77 60.68 68.01 45.58 33.42
Problem-As-Query Retrieval 4.97 69.51 63.14 53.65 60.93 65.22 75.06 31.06
CEDAR 5.82 72.44 62.63 53.71 61.61 61.68 75.06 32.67
RepoCoder 6.48 73.78 62.37 52.83 60.81 67.27 71.75 34.16
PERC 6.61 76.04 63.78 54.21 64.10 69.81 76.49 34.35
Table 1: Pass@1 scores of GPT-3.5-Turbo-16k augmented with different strategies for retrieving few-shot examples from Python source pools across CodeContests, HumanEval, and MultiPL-E benchmarks. Boldface indicates the best values while underline indicates the second-highest accuracy.
 Benchmark CodeContests HumanEval MultiPL-E
Target PL Python Python Rust Ruby Lua Julia Go R
Data Availability High-resource High-resource High-resource $\xleftrightarrow{\hskip 28.45274pt\text{}\hskip 28.45274pt}$ Underrepresented
Few-shot Prompting Method Pass@1 (%)
w/o Examples 4.97 87.87 81.22 78.88 66.96 61.51 52.73 49.07
Random Selection 5.58 86.59 81.35 79.50 73.98 68.36 47.01 53.42
Problem-As-Query Retrieval 7.64 87.07 80.58 80.12 72.92 65.97 68.83 52.80
CEDAR 8.18 88.17 81.15 80.75 74.04 69.37 71.43 53.42
RepoCoder 7.33 86.46 81.54 81.99 75.65 67.42 70.71 55.28
PERC 8.48 88.17 82.95 83.85 75.22 70.69 71.69 57.14
Table 2: Pass@1 scores of GPT-4o-mini augmented with different strategies for retrieving few-shot examples from Python source pools across CodeContests, HumanEval, and MultiPL-E benchmarks.
4 PERC
PERC retrieves relevant examples using algorithmic plans in pseudocode. These plans capture high-level logic while minimizing cross-lingual lexical differences, thereby supporting accurate code generation.
As depicted in Figure 2, the workflow of PERC consists of two key steps. First, it drafts a plan for the given problem to form an expanded query, which is used to retrieve examples that were projected to plan space in indexing time. Then, the retrieved examples and their plans are integrated into a reasoning chain to generate a revised plan and the final code.
4.1 Plan-As-Query Example Retrieval
A key contribution of PERC is the use of algorithmic plans written in pseudocode, for effective retrieval. Specifically, an LLM generates pseudocode $\hat{p}$ for the retrieval pool $\mathcal{P}$ offline, and $\hat{p}_{q}$ for $t_{q}$ at inference time. Then, the query is expanded with $\hat{p}_{q}$ as follows:
$E_{\textrm{PERC}}=\mathrm{topk}_{(t,\hat{p},c)}\ \mathrm{sim}(\psi(t_{q};\hat{p}_{q}),\psi(t;\hat{p})),$ (4)
where the in-context example for $\hat{p}_{q}$ is provided in Appendix C. As illustrated in Figure 1, Eq (4) avoids surface-level details by projecting code $c$ into plans. This is in contrast to Eq (3), which exposes the retriever to such distractions, especially when $\hat{c}$ and $c$ use different PLs.
4.2 Code Generation with Examples
We use generated pseudocode as intermediate reasoning steps for code generation Jiang et al. (2024). Each few-shot example in our prompt consists of a triplet $(t,\hat{p},c)$, where text description $t$, generated pseudocode $\hat{p}$, and code $c$ guide the LLM to utilize pseudocode in its reasoning chain:
$\texttt{prompt}=[[t;\hat{p};c]_{(t,\hat{p},c)\in E_{\textrm{PERC}}};t_{q}].$ (5)
When the target programming language differs from that of example code $c$, we replace $c$ with generated code $\hat{c}$ in the target language, where the LLM generates $\hat{c}$ using the in-context example shown in Appendix C:
$\texttt{prompt}=[[t;\hat{p};\hat{c}]_{(t,\hat{p},c)\in E_{\textrm{PERC}}};t_{q}].$ (6)
5 Experiments
5.1 Experimental Setup
Experiments were conducted using GPT-3.5-Turbo-16k and GPT-4o-mini as the backbone LLMs. Other implementation details regarding the embedding-based retrieval and hyperparameter configuration for code generation can be found in Appendix A.
Metrics
We evaluated the performance of PERC using the widely used Pass@1 metric Chen et al. (2021), which is an unbiased estimator of the model’s chance of producing a correct code sample in a single attempt.
Baselines
We compared our method against several established baselines to highlight the effectiveness of PERC: 1) w/o Examples generates code directly without using few-shot examples, 2) Random Selection uses a randomly chosen, then fixed set of examples, 3) Problem-As-Query Retrieval retrieves examples based on problem-problem similarity, 4) CEDAR Nashid et al. (2023) uses code-code similarity, and 5) RepoCoder Zhang et al. (2023) expands the query with predicted code.
Datasets
For evaluation, we used CodeContests Li et al. (2022), HumanEval Chen et al. (2021), and MultiPL-E (HumanEval; Cassano et al., 2023) benchmarks. For CodeContests, we used the train split as the example pool, while MBPP Austin et al. (2021) benchmark was used for the other two. Throughout the benchmarks, we used Python—a high-resource PL—as the source. We used Python as the primary target PL since it is the only officially supported language in both CodeContests and HumanEval benchmarks. For additional target PLs, we selected languages based on the frequency classes (NICHE, LOW, MEDIUM) established in MultiPL-E. We randomly chose two PLs from each class: Ruby and Go (MEDIUM), Rust and R (LOW), and Lua and Julia (NICHE).
5.2 RAG from Same PL Pool: CodeContests and HumanEval
Table 1 shows that PERC outperforms all baselines, achieving Pass@1 scores of 6.61% and 76.04% on CodeContests and HumanEval, respectively, using the GPT-3.5-Turbo-16k model. Similarly, the results for GPT-4o-mini in Table 2 show Pass@1 scores of 8.48% and 88.17%, demonstrating consistently high performance across benchmarks. This supports that retrieval based on algorithmic plans better captures the logic of the code and allows to surface more effective demonstrations in top-$k$.
5.3 RAG from Cross-PL Pool: MultiPL-E
The results for each PL in MultiPL-E, presented in Tables 1 and 2, are sorted in descending order of code generation accuracy without examples. PLs with higher accuracy are considered to have higher data availability, while those with lower accuracy are regarded as underrepresented.
Using the GPT-3.5-Turbo-16k model, PERC achieves the best Pass@1 scores across all PLs, with notable results such as 69.81% for Ruby, 63.78% for Rust, and 64.10% for Lua, as shown in Table 1. The results for GPT-4o-mini, presented in Table 2, also emphasizes PERC’s effectiveness, showing consistent Pass@1 score improvements, including 83.85% for Ruby, 70.69% for Julila, and 57.14% for R.
By effectively transferring knowledge from high-resource PLs, PERC demonstrated improved code generation accuracy for different PLs, showing its ability to bridge knowledge gaps across PLs with different data availability. In contrast, state-of-the-art approaches RepoCoder and CEDAR struggled with code generation. This limitation stemmed from their reliance on code-based retrieval, where modality differences introduced noise and hindered the identification of algorithmically relevant code. [^1] One may consider a cost-exhaustive approach of translating all the code in the pool to target (underrepresented) PLs, which incurs $\mathcal{O}(|\mathcal{T}||\mathcal{P}|)$ cost where $\mathcal{T}$ is the set of target PLs to handle, whereas PERC only requires $\mathcal{O}(|\mathcal{P}|)$.
6 Analysis and Discussion
Open-Source LLM as a Backbone
Table 3 shows consistent accuracy improvements with the open-source model Llama-3.1-8B-Instruct. PERC outperformed the baselines and demonstrated effective performance across PLs like Ruby, Lua, and R in the MultiPL-E benchmark. This highlights the improvements with PERC generalizes well to smaller, public backbone models.
 Benchmark MultiPL-E
Target PL Ruby Lua R
Method Pass@1 (%)
w/o Examples 46.09 39.63 18.82
Random Selection 45.34 38.70 22.36
RepoCoder 46.21 41.06 18.94
PERC 47.33 44.22 23.66
Table 3: Pass@1 scores of Llama-3.1-8B-Instruct augmented with different strategies for retrieving few-shot examples on Python source pool, on MultiPL-E benchmarks.
 Benchmark CodeContests MultiPL-E
Cand. PL C++ Java C++ Java
Target PL Python Python Lua Lua
RepoCoder 05.45 05.94 58.88 58.01
PERC 06.61 06.06 64.60 64.60
Table 4: Pass@1 scores of PERC and RepoCoder when using C++ and Java candidates in the CodeContests and MultiPL-E Lua benchmarks.
 Cand. PL Python Python/C++ Python/C++/Java
RepoCoder 6.48 5.88 4.85
PERC 6.61 6.48 6.12
Table 5: Pass@1 scores on CodeContests as more examples from different PLs are added to the retrieval pool.
Figure 3: Source PL distribution of retrieved examples on CodeContests, under the Mixed PL Pool setting of Table 5. The target PL, Python, is highlighted in blue.
 Method Pass@$1$ (%)
PERC 64.10
- Converting to Target PL 55.28
Table 6: Pass@1 of PERC on MultiPL-E-Lua benchmark with and without code conversion to target PL.
 Method Pass@$1$ (%)
PERC w/ Converted Code 5.52
PERC w/ Gold Code 5.70
Table 7: Pass@1 difference when replacing generated target PL code with gold target PL code on subsets of example pools containing both C++ and Python code in CodeContets
Using C++ and Java as Source PLs
Table 4 shows that PERC is also effective when using source PLs other than Python, namely C++ and Java. These results demonstrate that selecting examples based on algorithmic plans, regardless of the source PLs, can enhance code generation accuracy. The accuracy improvements using C++ and Java code pool for all MultiPL-E benchmarks are detailed in Appendix B.
Mixed PL Pool
As shown in Table 5, when C++ and Java code were incrementally added to the retrieval pool, PERC maintained higher accuracy while RepoCoder suffered from severe performance degradation. Again, this showcases the adaptability and robustness of retrieving examples with plans, than with code.
As illustrated in Figure 3, PERC outperforms RepoCoder while retrieving less examples of the target PL, Python; this empirically supports that PERC retrieves useful examples in non-target PLs.
Converting Code to Target PL
As illustrated in Section 4.2, PERC uses the converted code $\hat{c}$ rather than the original code of the retrieved example $c$, if the source and target PLs do not match. Table 6 shows such conversion is crucial, as syntax, APIs, and other elements specific to the source PL may inadvertently influence the generation and lead to performance degradation. Additionally, as shown in Table 7, replacing the generated target PL code with the gold target PL code in CodeContests, [^2] The subset of examples with both C++ and Python GT code available was used for evaluation. resulted in a minimal Pass@1 difference. This indicates potential errors that can be introduced in conversion to target PL are negligible.
7 Conclusion
We presented PERC, a novel framework for code generation that utilizes algorithmic plans both at indexing and generation time to select more effective few-shot examples to guide LLM. PERC demonstrates notable improvements in Pass@1 on CodeContests and HumanEval benchmark which represent scenario where the target PL is of high-resource, and also on MultiPL-E benchmarks for targeting underrepresented PLs.
Limitations
While PERC brings notable performance improvements in code generation with few-shot prompting, even if the source and target PLs do not match, our observation in Section 6 is that PERC shows slight performance drop as more and more programming languages are added to the retrieval pool. With an ideal retrieval system suited to selecting the most effective examples, one should observe monotonically increasing performance as more candidates are added to the pool; we leave more investigation and improvements as future work.
Acknowledgments
The first author was supported by Samsung Electronics. This work was partly supported by Electronics and Telecommunications Research Institute (ETRI) grant funded by ICT R&D program of MSIT/IITP (2022-0-00995, Automated reliable source code generation from natural language descriptions), and the MSIT (Ministry of Science and ICT), Korea, under the ITRC (Information Technology Research Center) support program (IITP-2024-2020-0-01789) supervised by the IITP (Institute for Information & Communications Technology Planning & Evaluation).
Appendix A Implementation Details
Throughout the experiments, we used GPT-3.5-Turbo-16k and GPT-4o-mini-2024-07-18 as the base LLM; for decoding, we applied nucleus sampling with $p=0.95$ and sharpening with temperature $T=0.8$. In all retrieval experiments, we employed semantic search using embedding models for retrieval, namely MPNet Song et al. (2020) [^3] https://huggingface.co/sentence-transformers/all-mpnet-base-v2 as the encoder $\psi$. For a fair comparison, we also used this retriever for the RepoCoder baseline as well, which originally used a sparse bag-of-words model. The top-$k$ candidates were then selected based on MPNet-based embeddings using cosine similarity, ensuring consistency across all methods.
In all few-shot prompting-based code generation experiments, ICL was performed using an inference chain consisting of problem description, pseudocode, and code. To ensure fair comparison of retrieval methods, the code was replaced with code converted to the target PL, as done in PERC’s approach. Regarding the number of shots, we employed 3-shot prompting for all benchmarks, except CodeContests, where 1-shot prompting was necessary due to the model’s 16k token limit.
Our code generation implementation primarily relied on the LangChain library. [^4] https://github.com/langchain-ai/langchain To execute and evaluate the generated code, we used code from the CodeEval repository [^5] https://huggingface.co/spaces/evaluate-metric/code_eval hosted on Huggingface for running evaluating on CodeContests and HumanEval benchmarks. Additionally, we utilized code from the bigcode-evaluation-harness repository [^6] https://github.com/bigcode-project/bigcode-evaluation-harness to evaluate on the MultiPL-E benchmark suite.
Appendix B Using C++ and Java Candidates for MultiPL-E Benchmarks
 Benchmark MultiPL-E
Candidate PL C++ Java C++ Java C++ Java
Target PL Lua Lua Ruby Ruby R R
RepoCoder 58.88 58.01 68.51 65.28 30.43 33.91
PERC 64.60 64.60 71.18 66.15 31.68 34.66
Table 8: Experimental results comparing the Pass@1 of PERC and RepoCoder when using C++ and Java candidates in the MultiPL-E HumanEval-Lua, Ruby, and R benchmarks.
As shown in Table 8, even when the candidate programming languages are C++ and Java rather than Python, PERC outperforms RepoCoder in the MultiPL-E HumanEval-Lua, Ruby, and R benchmarks. These results demonstrate that selecting examples based on pseudocode, regardless of the candidate PLs, can improve code generation accuracy for underrepresented PLs by leveraging the knowledge from examples written in high-resource PLs.
Appendix C In-Context Learning Examples
We provide concrete examples of prompts used for in-context learning in Tables 9, 10, and 11. The final prompt is created by appending the target problem to 3-shot examples in the same format as shown in the tables.
  Problem Description
def count_bidirectional(test_list):
   """Write a function to count bidirectional tuple pairs.
[…]
Write a plan for the problem.
[…]
# Define a function to count bidirectional tuple pairs.
# Create a map to store pairs and their reverse occurrences.
# Initialize a counter for bidirectional pairs.
# Iterate over the input list of tuples.
   # Retrieve the two elements of the tuple.
   # Create both the original pair and the reverse pair.
   # Check if the reverse pair exists in the map.
      # If it does, increment the bidirectional pair count and decrement the occurrence of the reverse pair in the map.
   # If the reverse pair does not exist in the map, increment the occurrence of the original pair in the map.
# Return the count of bidirectional pairs.
Table 9: The in-context example used for predicting the algorithmic plan from the test problem for query expansion by PERC. The problem description example is from Austin et al. (2021).
  Problem Description
def count_bidirectional(test_list):
   """Write a function to count bidirectional tuple pairs.
[…]
def count_bidirectional(test_list):
   pair_count = {}
   bidirectional_count = 0
   for tup in test_list:
      a, b = tup[0], tup[1]
      original = (a, b)
      reverse = (b, a)
      if reverse in pair_count and pair_count[reverse] > 0:
         bidirectional_count += 1
         pair_count[reverse] -= 1
      else:
         pair_count[original] = pair_count.get(original, 0) + 1
   return bidirectional_count
[…]
Write a plan for the problem.
[…]
# Define a function to count bidirectional tuple pairs.
# Create a map to store pairs and their reverse occurrences.
# Initialize a counter for bidirectional pairs.
# Iterate over the input list of tuples.
   # Retrieve the two elements of the tuple.
   # Create both the original pair and the reverse pair.
   # Check if the reverse pair exists in the map.
      # If it does, increment the bidirectional pair count and decrement the occurrence of the reverse pair in the map.
   # If the reverse pair does not exist in the map, increment the occurrence of the original pair in the map.
# Return the count of bidirectional pairs.
Table 10: The in-context example used for converting code in source PL to algorithmic plan by PERC. The problem description example is from Austin et al. (2021).
  Problem Description
def count_bidirectional(test_list):
   """Write a function to count bidirectional tuple pairs.
[…]
def count_bidirectional(test_list):
   pair_count = {}
   bidirectional_count = 0
   for tup in test_list:
      a, b = tup[0], tup[1]
      original = (a, b)
      reverse = (b, a)
      if reverse in pair_count and pair_count[reverse] > 0:
         bidirectional_count += 1
         pair_count[reverse] -= 1
      else:
         pair_count[original] = pair_count.get(original, 0) + 1
   return bidirectional_count
[…]
# Define a function to count bidirectional tuple pairs.
# Create a map to store pairs and their reverse occurrences.
# Initialize a counter for bidirectional pairs.
# Iterate over the input list of tuples.
   # Retrieve the two elements of the tuple.
   # Create both the original pair and the reverse pair.
   # Check if the reverse pair exists in the map.
      # If it does, increment the bidirectional pair count and decrement the occurrence of the reverse pair in the map.
   # If the reverse pair does not exist in the map, increment the occurrence of the original pair in the map.
# Return the count of bidirectional pairs.
[…]
Convert to Lua code for the problem following the plan.
[…]
function count_bidirectional(test_list)
   local pair_count = {}
   local bidirectional_count = 0
   for _, tup in ipairs(test_list) do
      local a, b = tup[1], tup[2]
      local original = {a, b}
      local reverse = {b, a}
      if pair_count[reverse] and pair_count[reverse] > 0 then
         bidirectional_count = bidirectional_count + 1
         pair_count[reverse] = pair_count[reverse] - 1
      else
         pair_count[original] = (pair_count[original] or 0) + 1
      end
   end
   return bidirectional_count
end
Table 11: The in-context example used for converting code in source PL to target PL by PERC. The problem description example is from Austin et al. (2021).
Appendix D Retrieved Example and Generated Code with PERC
We provide concrete example of the few-shot examples retrieved by PERC and the resultant pseudocode and generated code on HumanEval, in which the target PL is high-resource (python), in Tables 12 and 13. We also provide examples from MultiPL-E-Lua, underrepresented target PL setting, in Tables 14 and 15.
  Problem Description $t_{q}$
from typing import List
def below_zero(operations: List[int]) -> bool:
   """ You’re given a list of deposit and withdrawal operations on a bank account that starts with
   zero balance. Your task is to detect if at any point the balance of account falls below zero, and
   at that point function should return True. Otherwise it should return False.
   > below_zero([1, 2, 3])
   False
   > below_zero([1, 2, -4, 5])
   True
[…]
Predicted Pseudocode $\hat{p}_{q}$
# Set the initial balance of the bank account to zero
# Iterate through each operation in the list
   # Update the balance of the bank account by adding the current operation
   # Check if the balance is less than zero
      # If it is, return True
# If the balance never falls below zero, return False
[…]
Retrieved Problem with Pseudocode $t;p$
Write a function to calculate the sum of the negative numbers of a given list of numbers.
# Initialize the sum of negative numbers to 0
# Iterate through each number in the list
   # Check if the number is negative
      # If it is negative, add it to the sum
# Return the sum of negative numbers
[…]
Retrieved Code $c$
def sum_negativenum(nums):
   sum_negativenum = list(filter(lambda nums:nums<0,nums))
   return sum(sum_negativenum)
Table 12: The example of pseudocode to extend the query for a target HumanEval problem #3, along with the problem, pseudocode, code, and converted code in the target PL of the retrieved candidate through PERC.
  Problem Description
from typing import List
def below_zero(operations: List[int]) -> bool:
   """ You’re given a list of deposit and withdrawal operations on a bank account that starts with
   zero balance. Your task is to detect if at any point the balance of account fallls below zero, and
   at that point function should return True. Otherwise it should return False.
   > below_zero([1, 2, 3])
   False
   > below_zero([1, 2, -4, 5])
   True
[…]
# Initialize a variable "balance" to 0
# Iterate through each operation in the given list
   # Add the operation to the balance
   # Check if the balance is less than 0
      # If it is, return True
[…]
def below_zero(operations: List[int]) -> bool:
   balance = 0
   for operation in operations:
      balance += operation
      if balance < 0:
         return True
   return False
Table 13: An example of pseudocode and code generated using few-shot prompting with PERC for HumanEval problem #3, based on examples from Table 12.
  Problem Description $t_{q}$
– You’re given a table of deposit and withdrawal operations on a bank account that starts with
– zero balance. Your task is to detect if at any point the balance of account falls below zero, and
– at that point function should return true. Otherwise it should return false.
– > below_zero(1, 2, 3)
– false
– > below_zero(1, 2, -4, 5)
– true
local function below_zero(operations)
[…]
# Set the initial balance of the bank account to zero
# Iterate through each operation in the list
   # Update the balance of the bank account by adding the current operation
   # Check if the balance is less than zero
      # If it is, return True
# If the balance never falls below zero, return False
[…]
Retrieved Problem with Pseudocode $t;p$
Write a python function to return the negative numbers in a list.
# Initialize an empty list to store the negative numbers
# Iterate through each element in the input list
   # Check if the current element is less than 0
      # If the element is negative, append it to the list of negative numbers
# Return the list of negative numbers
[…]
Retrieved Code $c$
def neg_nos(list1):
   out = []
   for num in list1:
      if num < 0:
         out.append(num)
   return out
[…]
Converted Code in Target PL $\hat{c}$
function getNegativeNumbers(numbers)
   local negativeNumbers = {}
   for _, num in ipairs(numbers) do
      if num < 0 then
         table.insert(negativeNumbers, num)
      end
   end
   return negativeNumbers
end
Table 14: The example of pseudocode to extend the query for a target HumanEval-Lua problem #3, along with the problem, pseudocode, code, and converted code in the target PL of the retrieved candidate through PERC.
  Problem Description
– You’re given a table of deposit and withdrawal operations on a bank account that starts with
– zero balance. Your task is to detect if at any point the balance of account fallls below zero, and
– at that point function should return true. Otherwise it should return false.
– > below_zero(1, 2, 3)
– false
– > below_zero(1, 2, -4, 5)
– true
local function below_zero(operations)
[…]
# Return the count of bidirectional pairs.
# Initialize the balance of the account to 0
# Iterate through each operation in the list
   # Add the operation to the balance
   # Check if the balance is less than 0
      # If it is, return true
# If the loop completes without returning true, return false
[…]
function below_zero(operations)
   local balance = 0
   for _, operation in ipairs(operations) do
      balance = balance + operation
      if balance < 0 then
         return true
      end
   end
   return false
end
Table 15: An example of pseudocode and code generated using few-shot prompting for the target HumanEval-Lua problem #3, based on PERC-retrieved examples from Table 14.